Styles 1. The expected life expectancy for a sample of 6 dog breeds was collected. Breed Life Expectancy Cockapoo Dachshund English Bulldog Pointer Malamute 18 12 10 13 A. Calculate the 1QR. Please note, you may want to refer back to the lesson 2 lab as the 5 number summary was calculated there for this dataset. B. Determine if there are any outliers using the IQR rule. If there are outliers, note their values.
Answer #1

A.

Minimum = MIN(E14:E19) 8
Maximum = MAX(E14:E19) 18
First quartile, Q1 = QUARTILE.EXC(E14:E19,1) 9.5
Third quartile, Q3 = QUARTILE.EXC(E14:E19,3) 14.25
IQR = Q3-Q1 4.75

B. Lower limit = Q1 -1.5IQR = 9.5 - 1.5*4.75 = 2.375

Upper limit = Q3 + 1.5IQR = 14.25 + 1.5*4.75 = 21.375

As all the value are above lower limit and below upper limit, we can say that there are no outliers.